Walk-thru summary, Oct. 25

Angela Bellavance (bellavan@rice.edu)
Tue, 26 Oct 1999 08:54:38 -0500 (CDT)

Shift Summary for walk-thru shift, 25/Oct/99

On Shift: Angela Bellavance

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: accel. shutdown

Accesses during shift: Ramberg to look for phototubes for D0

Overview of shift:
Mostly quiet.
Had trouble with the SDAQ not fresh-starting correctly - Ashkan notified.
Shut off the terminals that usually display run histograms (BobT and
Chambers are at 2000V (Hogan).
Gas checklist done at 13:30.
Red buttons on the panel next to the restrooms:
- high water level duplex sump pump SP-2
- CUH-100 low temperature in vestibule 170
- water detected under floor in room 110 (labelled as "usually on")
Also noticed the control room is getting chilly.

Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke